A hydrogen atom be made up of a single proton and a single electron. The proton has a charge of +e and the electron has +e . In the “ground state” of the atom, the electron orbits the proton at most probable distance of 5.39 × 10−11 m. Determine the electric force on the electron due to the proton.
1
Expert's answer
2020-05-11T20:11:11-0400

According to Newton's third law, there is no difference between the force on the electron due to the proton or the force on the proton due to the electron.

The electric force can be calculated with Coulomb's law equation:


"F=\frac{1}{4\pi\epsilon_0}\cdot\frac{q_1q_2}{r^2}=\frac{1}{4\pi\epsilon_0}\cdot\frac{e^2}{r^2}=\\
\space\\
=\frac{1}{4\pi\cdot8.85\cdot10^{-12}}\cdot\bigg(\frac{1.60\cdot10^{-19}}{5.39 \cdot10^{−11}}\bigg)^2=7.94\cdot10^{-8}\text{ N}."